how do you solve systems with subtitute's
can please solve a few of these example problems so I get an ideal of what I need to do.
1)2x+7y=3 x=1-4y
2)4c=3d+3 c=d-1
3)3x-5y=11 x-3y=1
4)3x-2y=12 x=2y=6

Answer by elima(1433)

You are already given what one variable =, so plug that into the equation to solve for the other variable.
1)2x+7y=3 x=1-4y
2(1-4y)+7y=3
2-8y+7y=3
2-y=3
y=3-2
y=1
2x+7(1)=3
2x+7=3
2x=3-7
2x=-4
x=-2
See if you can solve the others, if you have problems let me know.
=)